Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

The probability is 1/2 that a certain coin will turn...

Expert replies
by swerve » Fri Feb 02, 2018 6:14 am
The probability is 1/2 that a certain coin will turn up heads on any given toss. If the coin is to be tossed four times, what is the probability that on at least one of the tosses the coin will turn up tails?

A. 1/16
B. 1/8
C. 1/2
D. 7/8
E. 15/16

The OA is E.

Please, can any expert explain this PS question for me? I can't get the correct answer. I need your help. Thanks.
Join the discussion
Source: — Problem Solving |

by ErikaPrepScholar » Fri Feb 02, 2018 6:49 am
Solving directly for the probability that we will get tails at least once is somewhat complicated — we have to think about the probability of getting tails on one, two, three, or four tosses, as well as in which order the tails will appear. However, we can solve this problem easily by instead finding the probability that we will get heads every single toss, then subtracting from one. This gives us the probability that we will NOT get heads every single toss — in other words, we will get tails at least once.

The probability we will get heads on each toss is 1/2. Multiplying by 1/2 for each toss gives
$$\frac{1}{2}\cdot\frac{1}{2}\cdot\frac{1}{2}\cdot\frac{1}{2}=\frac{1}{16}$$
So the probability that we will get heads every single toss is 1/16.

Subtracting this from one gives
$$1-\frac{1}{16}=\frac{15}{16}$$
So the probability that we will get tails at least once is 15/16, or answer choice E.

When solving for the probability that something will occur, it's often easiest to subtract out the cases that won't occur. Not always the case, but useful to keep in your tool belt!
Image

Erika John - Content Manager/Lead Instructor
https://gmat.prepscholar.com/gmat/s/

Get tutoring from me or another PrepScholar GMAT expert: https://gmat.prepscholar.com/gmat/s/tutoring/

Learn about our exclusive savings for BTG members (up to 25% off) and our 5 day free trial

Check out our PrepScholar GMAT YouTube channel, and read our expert guides on the PrepScholar GMAT blog
Join the discussion

by Brent@GMATPrepNow » Fri Feb 02, 2018 7:43 am
swerve wrote:The probability is 1/2 that a certain coin will turn up heads on any given toss. If the coin is to be tossed four times, what is the probability that on at least one of the tosses the coin will turn up tails?

A. 1/16
B. 1/8
C. 1/2
D. 7/8
E. 15/16
When it comes to probability questions involving "at least," it's best to try using the complement.
That is, P(Event A happening) = 1 - P(Event A not happening)
So, here we get: P(getting at least 1 tails) = 1 - P(NOT getting at least 1 tails)
What does it mean to NOT get at least 1 tails? It means getting zero tails.
So, we can write: P(getting at least 1 tails) = 1 - P(getting zero tails)

Now let's calculate P(getting zero tails)
What needs to happen in order to get zero tails?
Well, we need heads on the 1st toss and heads on the 2nd toss and heads on the 3rd toss and heads on the 4th toss
We can write P(getting zero tails) = P(heads on 1st AND heads on 2nd AND heads on 3rd AND heads on 4th)
This means that P(getting zero tails) = P(heads on 1st) x P(heads on 2nd) x P(heads on 3rd) x P(heads on 4th)
Which means P(getting zero tails) = (1/2)x(1/2)x(1/2) x(1/2) = 1/16

We're now ready to answer the question.
P(getting at least 1 tails) = 1 - P(NOT getting at least 1 tails)
= 1 - 1/16
= 15/16
= E

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by Scott@TargetTestPrep » Fri Jul 12, 2019 5:57 pm
swerve wrote:The probability is 1/2 that a certain coin will turn up heads on any given toss. If the coin is to be tossed four times, what is the probability that on at least one of the tosses the coin will turn up tails?

A. 1/16
B. 1/8
C. 1/2
D. 7/8
E. 15/16

.
The only way no tails will turn up is all four tosses turn up heads, which has a probability of 1/2 x 1/2 x 1/2 x 1/2 = 1/16. Therefore, the probability that at least one toss will turn up tails is 1 - 1/16 = 15/16.

Answer: E

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion